OpenStudy (anonymous):

Consider integer solutions to the equation: \[x_1+x_2+x_3+x_4+x_5=10\] How many non-negative integer solutions are there such that \(x_1 = x_2\)?

OpenStudy (anonymous):

@hartnn :D

ganeshie8 (ganeshie8):

say, \(x_1 = x_2 = x\) \(2x + x_3 + x_4 + x_5 = 10\) \(x_3 + x_4 + x_5 = 2(5-x)\) \(x = 0,1,2,3,4\)

ganeshie8 (ganeshie8):

add all 5 cases

OpenStudy (anonymous):

Hmm... Why can't x be 5?

OpenStudy (anonymous):

If x=5, we have \(x_3=x_4=x_5 = 0\)

ganeshie8 (ganeshie8):

oh yah it can be 5 also thats trivial case x3 = x4=x5 = 0

OpenStudy (anonymous):

Hmmm \(x_3+x_4+x_5=2(5−x)\) When x=0: \(x_3+x_4+x_5=10\) No of solutions = C(12, 2) = 66 When x=1: \(x_3+x_4+x_5=8\) No of solutions = C(10, 2) = 45 When x=2: \(x_3+x_4+x_5=6\) No of solutions = C(8, 2) = 28 When x=3: \(x_3+x_4+x_5=4\) No of solutions = C(6, 2) = 15 When x=4: \(x_3+x_4+x_5=2\) No of solutions = C(4, 2) = 6 When x=5: \(x_3+x_4+x_5=0\) No of solutions = C(2, 2) = 1 Total number of solutions = 66+45+28+15+6+1=161?!

ganeshie8 (ganeshie8):

\(\binom{12}{2} + \binom{10}{2} +\binom{8}{2} +\binom{6}{2} +\binom{4}{2} + 1\)

OpenStudy (anonymous):

Thanks! Coming up next: Number of solutions again!! :D

ganeshie8 (ganeshie8):

hah sure, these are fun if we know stars and bars !!

OpenStudy (anonymous):

We know stars and bars :D